OpenStudy (df001):

How to find the exact value of Sin(pi/4)?

OpenStudy (df001):

\[\theta = \frac{ \pi }{ 4 }\]

OpenStudy (df001):

hello

OpenStudy (astrophysics):

Think of the special triangle  this is the triangle for pi/4 use your trig ratios

OpenStudy (df001):

Is that a special triangle?

OpenStudy (astrophysics):

Yes

OpenStudy (astrophysics):

From unit circle

OpenStudy (df001):

Is it, 1/rad2?

OpenStudy (astrophysics):

There's one for pi/4, pi/3 and pi/6 but you can also derive them using the unit circle or just remember it

OpenStudy (astrophysics):

Yup exactly

OpenStudy (df001):

Sin=opp/hyp, 1/rad2

OpenStudy (df001):

okei thankss

OpenStudy (astrophysics):

Np

OpenStudy (df001):

excuse me, if it is asking for the exact value and 1/rad2 has a radical on the denominator, do I have to get rid of it by multiplying the numerator and denominator by radical 2?

OpenStudy (astrophysics):

You can it shouldn't really matter

OpenStudy (astrophysics):

\[\frac{ 1 }{ \sqrt{2} } = \frac{ \sqrt{2} }{ 2 }\]

OpenStudy (df001):

Thanks